# HuiduRep: A Robust Self-Supervised Framework for Learning Neural Representations from Extracellular Recordings

### v2

@@ Abstract
Extracellular recordings are transient voltage fluctuations in the vicinity of neurons, serving as a fundamental modality in neuroscience for decoding brain activity at single-neuron resolution. Spike sorting, the process of attributing each detected spike to its corresponding neuron, is a pivotal step in brain sensing pipelines. However, it remains challenging under low signal-to-noise ratio (SNR), electrode drift, and cross-session variability. In this paper, we propose HuiduRep, a robust self-supervised representation learning framework that extracts discriminative and generalizable features from extracellular recordings. By integrating contrastive learning with a denoising autoencoder, HuiduRep learns latent representations robust to noise and drift. With HuiduRep, we develop a spike sorting pipeline that clusters spike representations without ground truth labels. Experiments on hybrid and real-world datasets demonstrate that HuiduRep achieves strong robustness. Furthermore, the pipeline outperforms state-of-the-art tools such as KiloSort4 and MountainSort5. These findings demonstrate the potential of self-supervised spike representation learning as a foundational tool for robust and generalizable processing of extracellular recordings.
@@ Introduction
Neuroscientists frequently record extracellular action potentials, known as spikes, to monitor brain activity at single-cell resolution. These spikes, the extracellular voltage deflections from individual neurons, are considered the “fingerprints” of single-cell activity. By analyzing spike trains, sequences of temporally ordered spike times, researchers can infer neuronal coding and dynamics with millisecond precision (Bod et al. 2022).
However, each electrode often captures spikes from many nearby neurons, so it is crucial to sort or cluster spikes by their source (Dallal et al. 2016; Banga et al. 2022). Spike sorting is the process of assigning each detected spike waveform to its originating neuron. (Guzman et al. 2021) In practice, spike sorting is treated as a clustering problem on waveform features, often following initial steps of filtering and spike detection. (Souza et al. 2019) It is a foundational step in electrophysiology that enables single-unit analysis and studies of neuronal function (Rey, Pedreira, and Quian Quiroga 2015).
In classical spike sorting pipelines, data are first preprocessed, typically filtered and normalized. Spikes are then detected, typically via threshold crossings or template matching. Subsequently, features such as waveform principal components or wavelet coefficients are extracted. The resulting feature vectors are then clustered using methods like k-means, Gaussian Mixture Model (GMM), or density-based algorithms to identify putative single units. Early automated sorters such as KlustaKwik (Kadir, Goodman, and Harris 2013) often required extensive manual curation due to imperfect clustering. More recent frameworks like MountainSort (Chung et al. 2017) and KiloSort (Vishnubhotla et al. 2023) have improved throughput. For instance, MountainSort introduced an automatic clustering approach with accuracy comparable to or exceeding manual sorting. Likewise, KiloSort4 (Pachitariu et al. 2024) uses template matching and deconvolution to scale sorting to hundreds of channels with high accuracy. These tools represent the state-of-the-art in spike sorting, but they still rely on conventional clustering paradigms and presuppose stable, high-quality signals.
Despite recent advances, spike sorting remains challenging under realistic conditions. Low SNR signals make spikes hard to detect or distinguish. Nearby neurons often produce overlapping or morphologically similar waveforms, leading to ”compound” spikes that violate the assumption of one spike per neuron. Electrode drift, slow movement of neurons relative to the probe, causes spike waveforms to change over time, violating the stationarity assumption. Electrode drift has been identified as a major contributor to sorting errors, and correcting for drift substantially improves sorting performance. Spatial overlap of neurons also complicates sorting: dense, high–channel-count probes produce many overlapping electrical fields, worsening the ”collision” problem.
@@
We design a complete pipeline for spike sorting requires no ground truth labels and supports high-density probes.
We evaluate our method on datasets from distinct neural structures, demonstrating its robustness, and show that it outperforms state-of-the-art sorters.
@@
The overall architecture of HuiduRep is illustrated in Figure 1. Inspired by BYOL (Grill et al. 2020), our framework also consists of two main branches: an online network and a target network. The target network is updated via a momentum update based on the online network’s parameters, which are frozen during training.
The key difference lies in the introduction of a DAE within the online network, which is designed to reconstruct the original signals from the augmented views generated by the view generation module. This DAE serves as an auxiliary module to guide representation learning. Moreover, we replace the original ResNet encoder (He et al. 2015) in BYOL with a Transformer encoder (Vaswani et al. 2023). Before feeding the input views into the encoder, we also apply cross-channel convolution to better capture the characteristics of spike waveforms. During training, only View 1 is fed into the DAE branch, while View 2 does not participate in the denoising task.
Furthermore, the contrastive learning branch adapts the MoCo v3 style (Chen, Xie, and He 2021), where representations from positive pairs (query and key) and in-batch negative samples are compared. For contrastive learning, we adopt the InfoNCE loss (van den Oord, Li, and Vinyals 2019), while for denoising, we employ the mean squared error (MSE) loss:
Here denotes the query vector output by the prediction head of the online network, represents the positive key generated by the target network for the same sample and refers to the negative keys, which are the outputs of other samples in the same batch passed through the target network. is a temperature hyper-parameter. For MSE loss, is the embedded feature obtained from the original input, while is the reconstruction produced by the DAE. We apply a standard MSE loss to measure the reconstruction quality.(Wu et al. 2018) for -normalized and . The overall loss function of the model is a weighted sum of the denoising loss and the contrastive loss.
To generate input views, several augmentation strategies are employed to the original spike waveforms. These include: (1) Voltage and temporal jittering, which introduces small perturbations in both voltage amplitude and timing; (2) Channel cropping, where a random subset of channels is selected to create partial views of the original waveforms; (3) Collision, where noisy spikes are overlapped onto the original waveforms to simulate spike collisions; and (4) Noise, where temporally correlated noise is added to the waveforms to generate noised views. This Noise method is employed only for generating View 1, enhancing the robustness and performance of the DAE. The detailed view augmentation strategy is provided in the supplementary material.
@@
Our pipeline is modular, meaning that each component can be replaced by alternative methods. For example, the threshold-based detection module can be substituted with more accurate detection algorithms. In the following experiments, we demonstrate that even when using a threshold-based detection module with relatively low accuracy, our pipeline still outperforms the state-of-the-art and most widely used models such as Kilosort4.
@@
In this section, the datasets used for training and evaluating our model are presented, as well as their characteristics.
@@
DY016 and DY009 recordings are selected from the datasets released by IBL to train and evaluate HuiduRep. Both recordings were recorded from the hippocampal CA1 region and anatomically adjacent areas. Similar to the processing in CEED (Vishnubhotla et al. 2023), we used KiloSort2.5 (Pachitariu, Sridhar, and Stringer 2023) to preprocess the recordings and extracted a subset of spike units labeled as good according to IBL’s quality metrics (Banga et al. 2022) to construct our dataset. For every unit, we randomly selected 1,200 spikes for training and 200 spikes for evaluation. For each spike, we extracted a waveform with 121 samples across 21 channels, centered on the channel with the highest peak amplitude.
@@
HYBRID_JANELIA is a synthetic extracellular recording dataset with ground truth spike labels, designed to evaluate spike sorting algorithms. It was generated by using the Kilosort2 eMouse (Pachitariu, Sridhar, and Stringer 2023). The simulation includes a sinusoidal drift pattern with amplitude and 2 cycles over 1,200 seconds, as well as waveform templates from high-resolution electrode recordings.
We evaluated model performance on both the static and drift recordings of this dataset. To ensure a fair comparison, we reported results only on spike units with SNR greater than 3 for all models.
@@
In this section, we will introduce the key experimental procedures, including hyperparameter search, performance evaluation, and ablation studies.
@@
As shown in Table 1, the best overall performance was achieved when , with the highest ARI score and the highest max value. Notably, both very low () and high values () led to decreased performance, indicating that a moderate contribution of the denoising branch is essential for improving robustness and the overall performance of HuiduRep.
@@
All models under different settings were trained for 300 epochs with a batch size of 4096 and a fixed random seed on a server with a single NVIDIA L40s GPU and CUDA 12.4. A local evaluation server with a single NVIDIA RTX 5080 GPU and CUDA 12.8 is used to perform all experiments. A complete list of training hyperparameters is provided in the supplementary material.
Performance Evaluation
@@
To evaluate the performance of the HuiduRep Pipeline in real-world spike sorting tasks, two publicly available datasets, Hybrid Janelia and Paired MEA64c Yger, are selected as test sets. Multiple spike sorting tools, including Kilosort series (Pachitariu, Sridhar, and Stringer 2023) and MountainSort series (Chung et al. 2017), were evaluated. The performance of Kilosort4 and MountainSort5 was evaluated on our local evaluation server. The results for SimSort were cited from its original publication (Zhang et al. 2025), while the performance data for the remaining methods were obtained from the results provided by SpikeForest (Magland et al. 2020).
We recorded three metrics: accuracy (Acc), precision, and recall of different models across various test sets. Moreover, we adopted the SpikeForest definitions for computing these metrics, which slightly differ from the conventional calculation methods. The accuracy balances precision and recall, and it is similar to the F1-score. These metrics are computed based on the following quantities: : The number of ground-truth events that were missed by the sorter; : The number of ground-truth events that were correctly matched by the sorter; : The number of events detected by the sorter that do not correspond to any ground-truth event. Based on these definitions, the metrics are calculated as:
| Precision | |||
@@
| Dataset | without DAE | with DAE | ||||
| Distance | Silhouette Score | ARI | Distance | Silhouette Score | ARI | |
| IBL Test Dataset | ||||||
| Paired MEA64C 1 | N/A | N/A | ||||
| Paired MEA64C 2 | N/A | N/A | ||||
| Hybrid Janelia 1 | ||||||
| Hybrid Janelia 2 | ||||||
| Hybrid Janelia 3 | ||||||

### v3

@@ Abstract
Extracellular recordings are transient voltage fluctuations in the vicinity of neurons, serving as a fundamental modality in neuroscience for decoding brain activity at single-neuron resolution. Spike sorting, the process of attributing each detected spike to its corresponding neuron, is a pivotal step in brain sensing pipelines. However, it remains challenging under low signal-to-noise ratio (SNR), electrode drift and cross-session variability. In this paper, we propose HuiduRep, a robust self-supervised representation learning framework that extracts discriminative and generalizable features from extracellular recordings. By integrating contrastive learning with a denoising autoencoder, HuiduRep learns latent representations that are robust to noise and drift. With HuiduRep, we develop a spike sorting pipeline that clusters spike representations without ground truth labels. Experiments on hybrid and real-world datasets demonstrate that HuiduRep achieves strong robustness. Furthermore, the pipeline outperforms state-of-the-art tools such as KiloSort4 and MountainSort5. These findings demonstrate the potential of self-supervised spike representation learning as a foundational tool for robust and generalizable processing of extracellular recordings.
@@ Introduction
Neuroscientists frequently record extracellular action potentials, known as spikes, to monitor brain activity at single-cell resolution. These spikes, the extracellular voltage deflections from individual neurons, are considered the “fingerprints” of single-cell activity. By analyzing spike trains, which are sequences of temporally ordered spike times, researchers can infer neuronal coding and dynamics with millisecond precision (Bod et al. 2022).
However, each electrode often captures spikes from many nearby neurons, so it is crucial to sort or cluster spikes by their source (Dallal et al. 2016; Banga et al. 2022). Spike sorting is the process of assigning each detected spike waveform to its originating neuron (Guzman et al. 2021). In practice, spike sorting is treated as a clustering problem on waveform features, often following the initial steps of filtering and spike detection (Souza et al. 2019). It is a foundational step in electrophysiology that enables single-unit analysis and studies of neuronal function (Rey et al. 2015).
In classical spike sorting pipelines, data are first preprocessed, typically filtered and normalized. Spikes are then detected, typically via threshold crossings or template matching. Subsequently, features such as waveform principal components or wavelet coefficients are extracted. The resulting feature vectors are then clustered using methods like k-means, Gaussian Mixture Model (GMM) or density-based algorithms to identify putative single units. Early automated sorters such as KlustaKwik (Kadir et al. 2013) often required extensive manual curation due to imperfect clustering. More recent frameworks like MountainSort (Chung et al. 2017) and KiloSort (Vishnubhotla et al. 2023) have improved throughput. For instance, MountainSort introduced an automatic clustering approach with accuracy comparable to, or exceeding manual sorting. Likewise, KiloSort4 (Pachitariu et al. 2024) uses template matching and deconvolution to scale sorting to hundreds of channels with high accuracy. These tools represent the state-of-the-art in spike sorting, but they still rely on conventional clustering paradigms and presuppose stable, high-quality signals.
Despite recent advances, spike sorting remains challenging under realistic conditions. Low SNR signals make spikes difficult to detect or distinguish. Nearby neurons often produce overlapping or morphologically similar waveforms, leading to ”compound” spikes that violate the assumption of one spike per neuron. Electrode drift, slow movement of neurons relative to the probe, causes spike waveforms to change over time, violating the stationarity assumption. Electrode drift has been identified as a major contributor to sorting errors, and correcting for drift substantially improves sorting performance. Spatial overlap of neurons also complicates sorting: dense, high–channel-count probes produce many overlapping electrical fields, worsening the ”collision” problem.
@@
We design a complete pipeline for spike sorting that requires no ground truth labels and supports high-density probes.
We evaluate our method on datasets from distinct neural structures, demonstrating its robustness, and we show that it outperforms state-of-the-art sorters.
Code and supplementary materials are available at https://github.com/IgarashiAkatuki/HuiduRep
@@
In this section, we introduce the overall architecture of HuiduRep as well as the pipeline based on HuiduRep.
@@
The overall architecture of HuiduRep is illustrated in Figure 1. Inspired by BYOL (Grill et al. 2020) and MoCo-v3, our framework also consists of two main branches: an online network and a target network. The target network, which is frozen during training, is updated via a momentum update based on the online network’s parameters.
The key difference lies in the introduction of a DAE within the online network, which is designed to reconstruct the original spike representations from the augmented views generated by the view generation module. This DAE serves as an auxiliary module to guide representation learning. Moreover, we replace the original ResNet encoder (He et al. 2015) with a Transformer encoder (Vaswani et al. 2023). Before feeding the input views into the encoder, we also apply cross-channel convolution to better capture the characteristics of spike waveforms. During training, only View 1 is fed into the DAE branch, while View 2 does not participate in the denoising task.
Furthermore, the contrastive learning branch adapts the MoCo-v3 style (Chen et al. 2021), where representations from positive pairs (query and key) and in-batch negative samples are compared. For contrastive learning, we adopt the InfoNCE loss (van den Oord et al. 2019), while for denoising, we employ the mean squared error (MSE) loss:
Here denotes the query vector output by the prediction head of the online network, represents the positive key generated by the target network for the same sample and refers to the negative keys, which are the outputs of other samples in the same batch passed through the target network. is a temperature hyper-parameter (Wu et al. 2018) for -normalized and . For MSE loss, is the embedded feature obtained from the original input, while is the reconstruction produced by the DAE. We apply a standard MSE loss to measure the reconstruction quality. The overall loss function of the model is a weighted sum of the denoising loss and the contrastive loss.
Inspired by CEED, several augmentation strategies are employed to the original spike waveforms to generate input views. These include: (1) Voltage and temporal jittering, which introduces small perturbations in both voltage amplitude and timing; (2) Channel cropping, where a random subset of channels is selected to create partial views of the original waveforms; (3) Collision, where noisy spikes are overlapped onto the original waveforms to simulate spike collisions; and (4) Noise, where temporally correlated noise is added to the waveforms to generate noised views. This Noise method is employed only for generating View 1, enhancing the robustness and performance of the DAE. The detailed view augmentation strategy is provided in the supplementary material.
@@
Our pipeline is modular, meaning that each component can be replaced by alternative methods. For example, the threshold-based detection module can be substituted with more accurate detection algorithms. In the following experiments, we demonstrate that even when using a threshold-based detection module with relatively low accuracy, our pipeline still outperforms the state-of-the-art and most widely used methods such as KiloSort4.
@@
DY016 and DY009 recordings are selected from the datasets released by IBL to train and evaluate HuiduRep. Both recordings were recorded from the hippocampal CA1 region and anatomically adjacent areas. Similar to the processing in CEED (Vishnubhotla et al. 2023), we used KiloSort2.5 (Pachitariu et al. 2023) to preprocess the recordings and extracted a subset of spike units labeled as good according to IBL’s quality metrics (Banga et al. 2022) to construct our dataset. For every unit, we randomly selected 1,200 spikes for training and 200 spikes for evaluation. For each spike, we extracted a waveform with 121 samples across 21 channels, centered on the channel with the highest peak amplitude.
@@
HYBRID_JANELIA is a synthetic extracellular recording dataset with ground truth spike labels designed to evaluate spike sorting algorithms. It was generated by using the KiloSort2 eMouse (Pachitariu et al. 2023). The simulation includes a sinusoidal drift pattern with amplitude and 2 cycles over 1,200 seconds, as well as waveform templates from high-resolution electrode recordings.
We evaluated model performance on both the static and drift recordings of this dataset. To ensure a fair comparison, we reported results only on spike units with the SNR greater than 3 for all models.
@@
As shown in Table 1, the best overall performance was achieved when , with the highest ARI score and the highest maximum value. Notably, both very low () and high values () led to decreased performance, indicating that a moderate contribution of the denoising branch is essential for improving robustness and the overall performance of HuiduRep.
@@
All models under different settings were trained for 300 epochs with a batch size of 4096 and a fixed random seed on a server with a single NVIDIA L40s 48G GPU and CUDA 12.4. A local evaluation server with a single NVIDIA RTX 5080 16G GPU and CUDA 12.8 is used to perform all experiments. A complete list of training hyperparameters is provided in the supplementary material.
Performance Evaluation
To evaluate the performance of the HuiduRep Pipeline in real-world spike sorting tasks, two publicly available datasets, Hybrid Janelia and Paired MEA64c Yger, are selected as test sets. Multiple spike sorting tools, including KiloSort series (Pachitariu et al. 2023) and MountainSort series (Chung et al. 2017), were evaluated. The performance of KiloSort4 and MountainSort5 was evaluated on our local evaluation server. The results for SimSort were cited from its original publication (Zhang et al. 2025), while the performance data for the remaining methods were obtained from the results provided by SpikeForest (Magland et al. 2020).
We recorded three metrics: accuracy, precision, and recall of different models across various test sets. Moreover, we adopted the SpikeForest definitions for computing these metrics, which slightly differ from the conventional calculation methods. The accuracy balances precision and recall, and it is similar to the F1-score. These metrics are computed based on the following quantities: : The number of ground-truth events that were missed by the sorter; : The number of ground-truth events that were correctly matched by the sorter; : The number of events detected by the sorter that do not correspond to any ground-truth events. Based on these definitions, the metrics are calculated as:
| Precision | |||
| Dataset | without DAE | with DAE | ||||
| Distance | Silhouette Score | ARI | Distance | Silhouette Score | ARI | |
| IBL_Test_Dataset | ||||||
| Paired_MEA64C_Yger_1 | N/A | N/A | ||||
| Paired_MEA64C_Yger_2 | N/A | N/A | ||||
| Hybrid_Janelia_1 | ||||||
| Hybrid_Janelia_2 | ||||||
| Hybrid_Janelia_3 | ||||||
@@ Conclusion
Ethics Statement
This study uses only publicly available datasets, which were collected and shared in compliance with institutional and ethical guidelines as stated in the original publications.
Acknowledgments
We acknowledge Huidu, the ragdoll cat of the co-first authors, whose name inspired the title of the HuiduRep framework.